# JailBreakV-28K: A Benchmark for Assessing the Robustness of MultiModal Large Language Models against Jailbreak Attacks

### v2

@@ Abstract
With the rapid advancements in Multimodal Large Language Models (MLLMs), securing these models against malicious inputs while aligning them with human values has emerged as a critical challenge. In this paper, we investigate an important and unexplored question of whether techniques that successfully jailbreak Large Language Models (LLMs) can be equally effective in jailbreaking MLLMs. To explore this issue, we introduce JailBreakV-28282828111https://huggingface.co/datasets/JailbreakV-28K/JailBreakV-28k , a pioneering benchmark designed to assess the transferability of LLM jailbreak techniques to MLLMs, thereby evaluating the robustness of MLLMs against diverse jailbreak attacks. Utilizing a dataset of $2,000$ malicious queries that is also proposed in this paper, we generate $20,000$ text-based jailbreak prompts using advanced jailbreak attacks on LLMs, alongside $8,000$ image-based jailbreak inputs from recent MLLMs jailbreak attacks, our comprehensive dataset includes $28,000$ test cases across a spectrum of adversarial scenarios. Our evaluation of $10$ open-source MLLMs reveals a notably high Attack Success Rate (ASR) for attacks transferred from LLMs, highlighting a critical vulnerability in MLLMs that stems from their text-processing capabilities. Our findings underscore the urgent need for future research to address alignment vulnerabilities in MLLMs from both textual and visual inputs.
@@ 3.2 RedTeam-2K: A Comprehensive Malicious Query Dataset
We review some SotA jailbreak-related works for the first stages to find several datasets. Finally, we identified five datasets: LLM Jailbreak Study Liu et al. (2023b), AdvBench Zou et al. (2023), BeaverTails Ji et al. (2023), Question Set Shen et al. (2023), and hh-rlhf of Anthropic Bai et al. (2022). All other datasets have safety policy labels except for Advbench and hh-rlhf by manual annotation. Based on the OpenAI usage policies Ope (2024) and LLaMa-$2$ usage policy Meta AI (2024), we conducted a decomposition and extraction on datasets tagged with safety policy labels. This process involved identifying and extracting safety policy labels not encompassed within the existing OpenAI and LLaMa-$2$ usage policy. Subsequently, we reconstructed these extracted safety labels to formulate a distinct set of safety policy labels pertinent to our dataset, detailed in Table 1.
Category (Ours) Definition
Illegal Activity
Engage in unlawful conduct or activities(terrorism
hack, human trafficking, rape, etc.)
Violence Harming, threatening, or bullying individuals.
Hate Speech
Derogatory language directed at a specific group
usually incites harm.
Malware
Software, viruses, and scripts designed to harm or
exploit computer systems.
Physical Harm
High risk of bodily harm, weapons manufacturing
poisoning, suicide, and illegal manufacturing.
Economic Harm
Elevated risk of financial harm (including multi-
level marketing, gambling, etc.).
Fraud
Deceptive, misleading, and misinforming actions
for personal or financial gain.
Child Abuse Harm or neglect of a child physically or sexually.
Animal Abuse Illegal trade, poaching, or harm of animals.
Political Sensitivity
Political campaigning or lobbying and sensitive
topics in international politics.
Privacy Violation
Unauthorized access or exposure of personal
information.
Tailored Unlicensed Advice
Unauthorized/Unlicensed practice of professions
(including legal and financial).
Health Consultation
Offering guidance on treating or managing health
conditions and advising on medical issues.
Government Decision High risk government decision-making.
Unethical Behavior Legal, nonviolent, immoral behavior.
Bias
Racial bias, personal prejudice, and group-targeted
discrimination.
Table 1: Our Safe Policy Definition
@@ 3.2.3 LLM-based Generation
Large language models have demonstrated considerable efficacy in the generation of data Huang et al. (2023). After integrating all datasets, we aim to enrich each sparse safety policy by generating various harmful queries. We incorporated statements from OpenAI’s usage policy and the LLaMa-$2$ usage policy for prompt design within our prompts. Unlike the LLM-based generation method used in SafeBench Gong et al. (2023), in our methodology for prompting GPT, we intentionally included a variety of syntax, such as interrogative and imperative forms, along with multi-scenario descriptions of hypothetical real-life. This strategy was employed to enhance the diversity and realism of our dataset. Details of prompt engineering are shown in Figure 3.
@@ 4.2 Results and Analysis
The MLLMs inherit vulnerabilities from their LLM counterparts. We evaluated the ASR of these LLM transfer attacks on the LLM encoder of these MLLMs without images. Detailed in Table  5. The average ASR of these LLM encoders has achieved $68.7\%$. Additionally, these LLM transfer attacks initially generated against $8$ LLMs also achieved an average ASR of $64.4\%$ on these LLMs detailed in Table 6. These attacks still maintained high ASRs when applied to MLLMs, especially Template and Logic, detailed in Figure 6 in Appendix A. These patterns suggest that jailbreak prompts that successfully jailbreak LLMs can be effectively adapted and transferred to attack MLLMs, highlighting a significant vulnerability that spans across from LLMs to MLLMs and emphasizing the need for robust defense mechanisms that can adapt to the evolving nature of such LLM transfer attacks in MLLMs.
@@
LLM Template Persuade Logic Total
Vicuna-7b 92.8 91.8 94.6 92.7
Llama-7b 82.3 62.3 54.1 80.5
Llama-13b 80.8 77.2 47.3 80.1
Qwen1.5-7B 38.8 40.1 24.3 38.6
phi-2 50.2 54.7 58.1 50.6
Baichuan-7B 40.1 49.4 63.5 41.1
ChatGLM3-6B 60.5 66.3 62.2 61
Mixtral-8x7B-Instruct-v0.1 71.5 57.3 83.8 70.7

### v4

# JailBreakV: A Benchmark for Assessing the Robustness of MultiModal Large Language Models against Jailbreak Attacks

@@ Abstract
With the rapid advancements in Multimodal Large Language Models (MLLMs), securing these models against malicious inputs while aligning them with human values has emerged as a critical challenge. In this paper, we investigate an important and unexplored question of whether techniques that successfully jailbreak Large Language Models (LLMs) can be equally effective in jailbreaking MLLMs. To explore this issue, we introduce JailBreakV-28282828111https://huggingface.co/datasets/JailbreakV-28K/JailBreakV-28k , a pioneering benchmark designed to assess the transferability of LLM jailbreak techniques to MLLMs, thereby evaluating the robustness of MLLMs against diverse jailbreak attacks. Utilizing a dataset of $2,000$ malicious queries that are also proposed in this paper, we generate $20,000$ text-based jailbreak prompts using advanced jailbreak attacks on LLMs, alongside $8,000$ image-based jailbreak inputs from recent MLLMs jailbreak attacks, our comprehensive dataset includes $28,000$ test cases across a spectrum of adversarial scenarios. Our evaluation of $10$ open-source MLLMs reveals a notably high Attack Success Rate (ASR) for attacks transferred from LLMs, highlighting a critical vulnerability in MLLMs that stems from their text-processing capabilities. Our findings underscore the urgent need for future research to address alignment vulnerabilities in MLLMs from both textual and visual inputs.
@@ 3.2 RedTeam-2K: A Comprehensive Malicious Query Dataset
We review some SotA jailbreak-related works for the first stages to find several datasets. Finally, we identified five datasets: LLM Jailbreak Study Liu et al. (2023b), AdvBench Zou et al. (2023), BeaverTails Ji et al. (2023), Question Set Shen et al. (2023), and hh-rlhf of Anthropic Bai et al. (2022). All other datasets have safety policy labels except for Advbench and hh-rlhf by manual annotation. Based on the OpenAI usage policies Ope (2024) and LLaMa-$2$ usage policy Meta AI (2024), we conducted a decomposition and extraction on datasets tagged with safety policy labels. This process involved identifying and extracting safety policy labels not encompassed within the existing OpenAI and LLaMa-$2$ usage policy. Subsequently, we reconstructed these extracted safety labels to formulate a distinct set of safety policy labels pertinent to our dataset, detailed in Table 6 in Appendix A.
@@ 3.2.3 LLM-based Generation
Large language models have demonstrated considerable efficacy in the generation of data Huang et al. (2023). After integrating all datasets, we aim to enrich each sparse safety policy by generating various harmful queries. We incorporated statements from OpenAI’s usage policy and the LLaMa-$2$ usage policy for prompt design within our prompts. Unlike the LLM-based generation method used in SafeBench Gong et al. (2023), in our methodology for prompting GPT, we intentionally included a variety of syntax, such as interrogative and imperative forms, along with multi-scenario descriptions of hypothetical real-life. This strategy was employed to enhance the diversity and realism of our dataset. Details of prompt engineering are shown in Figure 4 in Appendix A.
@@ 4.2 Results and Analysis
The MLLMs inherit vulnerabilities from their LLM counterparts. We evaluated the ASR of these LLM transfer attacks on the LLM encoder of these MLLMs without images. Detailed in Table  4. The average ASR of these LLM encoders has achieved $68.7\%$. Additionally, these LLM transfer attacks initially generated against $8$ LLMs also achieved an average ASR of $64.4\%$ on these LLMs detailed in Table 8 in Appendix. These attacks still maintained high ASRs when applied to MLLMs, especially Template and Logic, detailed in Figure 6 in Appendix A. These patterns suggest that jailbreak prompts that successfully jailbreak LLMs can be effectively adapted and transferred to attack MLLMs, highlighting a significant vulnerability that spans across from LLMs to MLLMs and emphasizing the need for robust defense mechanisms that can adapt to the evolving nature of such LLM transfer attacks in MLLMs.
@@ 5 Conclusion
6 Acknowledgements
We would like to express our sincere gratitude to the anonymous reviewer(s) for their valuable feedback and constructive comments, which significantly contributed to the improvement of this paper.
7 Ethics Statement
This paper introduces a jailbreak benchmark designed to evaluate the robustness of multi-modal large Language Models (MLLMs) against jailbreak attacks. These attacks can potentially lead MLLMs to produce outputs that are misaligned with human preferences, intentions, or values. While we anticipate that this research, like previous work on jailbreaks, will not cause immediate harm, we aim to inspire the development of more effective defense strategies. This, in turn, will contribute to the creation of more robust, safe, and aligned LLMs over time.
The benchmark presented is based on jailbreak queries that have demonstrated high success rates against LLMs and strong efficacy against MLLMs. Consequently, these queries have the potential to prompt MLLMs to generate malicious outputs. To mitigate the risk of misuse, access to this benchmark for research or educational purposes requires prior permission via our designated data access link.
In the short term, this benchmark could facilitate a deeper understanding of vulnerabilities in MLLMs, potentially leading to immediate improvements in model safety and alignment. In the long term, the insights gained from this research are expected to drive the development of more resilient and ethically aligned LLMs, fostering a safer and more reliable interaction between humans and AI systems. By addressing these vulnerabilities proactively, we contribute to a future where AI technologies are better equipped to adhere to human values and intentions.
@@ Appendix A Appendix
Category (Ours) Definition
Illegal Activity
Engage in unlawful conduct or activities(terrorism
hack, human trafficking, rape, etc.)
Violence Harming, threatening, or bullying individuals.
Hate Speech
Derogatory language directed at a specific group
usually incites harm.
Malware
Software, viruses, and scripts designed to harm or
exploit computer systems.
Physical Harm
High risk of bodily harm, weapons manufacturing
poisoning, suicide, and illegal manufacturing.
Economic Harm
Elevated risk of financial harm (including multi-
level marketing, gambling, etc.).
Fraud
Deceptive, misleading, and misinforming actions
for personal or financial gain.
Child Abuse Harm or neglect of a child physically or sexually.
Animal Abuse Illegal trade, poaching, or harm of animals.
Political Sensitivity
Political campaigning or lobbying and sensitive
topics in international politics.
Privacy Violation
Unauthorized access or exposure of personal
information.
Tailored Unlicensed Advice
Unauthorized/Unlicensed practice of professions
(including legal and financial).
Health Consultation
Offering guidance on treating or managing health
conditions and advising on medical issues.
Government Decision High risk government decision-making.
Unethical Behavior Legal, nonviolent, immoral behavior.
Bias
Racial bias, personal prejudice, and group-targeted
discrimination.
Table 6: Our Safe Policy Definition
@@
LLM Template Persuade Logic Total
Vicuna-7b 92.8 91.8 94.6 92.7
Llama-7b 82.3 62.3 54.1 80.5
Llama-13b 80.8 77.2 47.3 80.1
Qwen1.5-7B 38.8 40.1 24.3 38.6
phi-2 50.2 54.7 58.1 50.6
Baichuan-7B 40.1 49.4 63.5 41.1
ChatGLM3-6B 60.5 66.3 62.2 61
Mixtral-8x7B-Instruct-v0.1 71.5 57.3 83.8 70.7